OpenStudy (anonymous):

solve the given equation to approximate the angle x. [0,pi/2) is the interval. Round to three decimal places: 75sin^2x - 34sinx+ 3 = 0

OpenStudy (anonymous):

(25sinx-3)(3sinx-1)=0 sinx=3/25 and sinx=1/3 after that, you should a calculator to find the answer.

OpenStudy (anonymous):

Oh wow alright that really helped! Thank you! :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

We put : \[t=\sin x\] Then, the equation becomes : \[75t^2-34t+3=0\] and : \[\Delta=34^2-4\times75\times 3=256=16^2>0\] So : \[t_1=\frac{34+16}{2\times75}=\frac13\\ t_2=\frac{34-16}{2\times75}=\frac{3}{25} \] So : \[\sin x=\frac13\Rightarrow x=0.34+2k\pi\text{ or }x=(\pi-.34)+2k\pi \\\sin x=\frac3{25}\Rightarrow x=0.12+2k\pi\text{ or }x=(\pi-.12)+2k\pi \]

OpenStudy (anonymous):

thank you :) @Noura11